How many solutions does -14x+7y=21
-8x+4y=12 have?

Found 2 solutions by stanbon, Hawksfan:
Answer by stanbon(75887)

How many solutions does
-14x+7y=21
-8x+4y=12 have?
-------
Modify::
y - 2x = 3
y - 2x = 3
------
Since the lines are the same, the system has an infinity of solutions
of the form:: (x,2x+3)
--------------------
Cheers,
Stan H.
---------------

Answer by Hawksfan(61)

-14x+7y=21
-8x+4y=12
If you multiply the top equation by -8 and the bottom equation by 14 so that the x's cancel each other out, you get...
-112x -56y = 168
112x +56y = -168
Surprise,surprise because when you add the equations together, they cancel each other totally out. This is a case when the graph of line 1 is the same as line 2. They are the same line. They are not parallel, but are the same line. There will be infinite amount of solutions.